a) A beam of 7.8-MeV α particles scatters from a gold foil of thickness 0.30 μm.
i) What fraction of the α particles is scattered between 1.5° and 2.5°?
ii) What is the ratio of α particles scattered through angles greater than 2° to the number
scattered through angles greater than 10°?
1
Expert's answer
2020-06-15T10:35:35-0400

I)


"f(1.5)-f(2.5)=\pi nt\left(\frac{Z_1Z_2e^2}{8\pi\epsilon_0K}\right)^2\\(\cot^2{0.5(1.5)}-\cot^2{0.5(2.5)})"

"f(1.5)-f(2.5)=\pi (5.9\cdot10^{-6})(0.3\cdot10^{28})\\\left(\frac{(2)(79)(1.6\cdot10^{-19})^2}{8\pi(8.85\cdot10^{-12})(7.8\cdot10^{6})(1.6\cdot10^{-19})}\right)^2\\(\cot^2{0.5(1.5)}-\cot^2{0.5(2.5)})"

"f(1.5)-f(2.5)=0.044"

II)


"\frac{f(2)}{f(10)}=\frac{\cot^2(0.5(2))}{\cot^2(0.5(10))}=25.1"